OpenStudy (anonymous):

A sequence is defined by the recursive function f(n + 1) = f(n). If f(3) = 9 , what is f(1) ?

OpenStudy (jango_in_dtown):

sure this is the question?? I dont see any recursion here... @destiny5

OpenStudy (jango_in_dtown):

and if the question is correct then 9=f(3)=f(2)=f(1) and hence f(1)=9

OpenStudy (freckles):

f(n+1)=f(n) <---is this right? f(n+1)=f(n) means term equals previous term which means all terms are equal